OpenStudy (anonymous):

Doug is a salesperson in a retail store and earns $75 per week, plus 15% of his weekly sales. If Doug makes $555 one week, what are his sales for that week? A. $3,200 B. $627 C. $566 D. $1,914

OpenStudy (ranga):

Let the weekly sales be S dollars that week. Doug earns $75 per week. He earns 15% commission on his weekly sales. So 75 + S * 0.15 = 555 solve for S

OpenStudy (anonymous):

How do i solve it?

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

75 + S * 0.15 = 555 To solve for S, isolate S. Subtract 75 from both sides: S * 0.15 = 555 - 75 = 480 divide both sides by 0.15 S = 480 / 0.15 = ?

OpenStudy (anonymous):

3,200 Yes! Thank You!

OpenStudy (ranga):

You are welcome.